OpenStudy (anonymous):

Helps: Prob :1 Use elimination to solve the system of equations given by -7x + 3y = 14 and 7x - 2y = -14 A: (-2, 0) B: (1, -2) C: (0, -2) D: (0, 2) Prob 2: Use elimination to solve the system of equations given by 10x - 40y = 60 and 15x + 40y = -10. A: (0, 1) B: (1, 1) C: (-2, -2) D: (2, -1) Notes: If you could teach me how to do it when you help me it would be grate, But trying to keep it fast

OpenStudy (anonymous):

there are 3 more just like it so i will try to get them right after someone help me.. so more Medals for every one xD

OpenStudy (anonymous):

Helps ^

OpenStudy (anonymous):

A and D

OpenStudy (anonymous):

Thanks but i need to learn it

OpenStudy (anonymous):

Math can to brake it down in to Small steps like supper small xD

OpenStudy (anonymous):

prob 2 10x - 40y = 60 and 15x + 40y = -10. add both equations you will get 25x=50 so x=50/25=2 now substitute x value in one of the equation 10(2)-40y=60 -40y=60-20 y=40/-40 y=-1

OpenStudy (anonymous):

so for Use elimination to solve the system of equations given by 3x - 2y = 12 and 3x + 5y = 33. I add 3x - 2y = 12 + 3x + 5y = 33

OpenStudy (anonymous):

prob 1 -7x + 3y = 14 and 7x - 2y = -14 add both equations you will get y=0 now substitute y value in one of the equation 7x-2(0)=-14 x=-14/7 x=-2

OpenStudy (anonymous):

Can you show how to add both equations

OpenStudy (anonymous):

-7x+3y=14 7x-2y=-14 ------------- 0+y=0

OpenStudy (anonymous):

3x - 2y = 12 and 3x + 5y = 33. in this case you should subtract 2nd equation from 1st equation

OpenStudy (anonymous):

err we need a whiteboard

OpenStudy (anonymous):

do you happen to know of one we can use?

OpenStudy (anonymous):

http://www.twiddla.com/580173

OpenStudy (anonymous):

can you show me on there